OpenStudy (anonymous):

Trisha needs 4 1/2 yards of fabric to do a project. If she already has 2/3 of a yard of fabric, what is the minimum amount of fabric she must buy to be able to complete her project? a.) 3 5/6 b.) 4 1/12 c.) 4 1/6 thanks in advance:)

OpenStudy (poopsiedoodle):

Trisha should just get the 4 1/2 yards and stop being a jew. That would make things less complicated.

OpenStudy (anonymous):

I'm sorry, "stop being a jew"?

OpenStudy (poopsiedoodle):

Anyway, I think it's C., but you might need someone to confirm it.

OpenStudy (poopsiedoodle):

Yes, Trisha should stop being a jew.

OpenStudy (aripotta):

4 1/2 - 2/3 = 9/2 - 2/3 = 27/6 - 4/6 = 23/6 = 3 5/6

OpenStudy (poopsiedoodle):

Ah. Well, I was wrong. :P

OpenStudy (anonymous):

thanks guys:)